Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for April 1999

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Friday, April 30, 1999

FOR DATES: From 7:00 a.m. Thursday, April 22, 1999 to 7:00 a.m. Friday, April 30, 1999.

The Thermal Destruction Facility (TDF) did not process any contaminated soil during the time frame of this report. The TDF was in a hot hold mode until late in the evening on Friday April 23,1999 at which time the final bottom ash analytical results were received from the project laboratory. The results indicated that the thermally treated soils were acceptable for backfill onsite and the contractor was instructed to begin demobilization of the TDF. The main valve on the natural gas line was closed, which caused the burners on the rotary kiln and secondary combustion chamber to shut off and the cool down of the TDF began. The cool down of the TDF was completed on Saturday, April 24, 1999. The contractor, OHM, is currently working 24 hours a day, 5 days a week demobilizing the TDF and all supporting equipment.

The final bottom ash analytical reports for the thermally treated soils processed April 18, 19, 20, and 21 for the contents of bins 1,3,5,& 7 were received. The analytical test results indicated that the thermally treated material met the project's organic criteria for backfill above elevation 547. Received flyash analytical results for the contents of bin C generated during the period April 13 to April 20. The analytical test results indicated thermally treated material met the project's criteria for backfill above elevation 553.

Site activities during the time frame of this report included cool down of the TDF, wastewater management, equipment decontamination, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. A nitrogen purge of both the natural gas line and the oxygen lines onsite were performed on separate days with the assistance of the vendors. The removal of the liquid oxygen storage tanks and vaporizer banks was completed. The decon and removal of the steam scrubbing unit was completed. The removal of the enclosure around the flyash storage building was completed. Flushing of the acid gas scrubber system and caustic storage tank was completed. The last bottom ash material was backfilled. The feed prep air handling system was removed. All screening equipment inside of the feed prep building was decontaminated. Contractor began cleaning the HDPE liner system in preparation for removal. The sub- contractor responsible for perimeter and off-site air monitoring stations retrieved the equipment and demobilized from site.

Related site activities included a weekly progress meeting followed by a monthly partnering session was held on Tuesday, April 27, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

Major activities for the next week includes the internal clean out of duct work and equipment using a vacuum truck, the removal of internal components of the acid gas scrubbing system and bag house unit, the collection and sampling of any materials generated during clean out of the TDF, and the shipment of equipment, debris and scrap metal off-site to various facilities.

PREPARED BY: Chad Thompson for Michael Ogden (USACE)


Thursday, April 22, 1999

FOR DATES: From 7:00 a.m. Wednesday, Apr. 14, 1999 to 7:00 a.m. Thursday, Apr. 22, 1999.

The Thermal Destruction Facility (TDF) was on-line for 169.3 hours and processed 5666.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 88.2 % of the time with an average hourly feed rate of 33.5 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully completed treating approximately 290,381.6 tons of contaminated soil, which amounts to 100% of the overall contaminated on-site soil.

Hazardous waste feed ended Thursday, April 22, 1999 at 0609 hours. The plant has shifted into a hot hold condition awaiting the completion of final ash analytical testing. It is expected the last analytical report will be received on Saturday, April 24,1999. If the remaining four bins meet the contract lay down criteria, the plant will be shut down and the project will go into the demobilization phase.

Hazardous waste feed was stopped on Wednesday, April 14, 1999 in order to investigate problems with the quench tower spray nozzles. The nozzles were cleaned, reassembled and reinstalled. Feed was off for approximately 10 hours. Monthly hazardous waste feed interlock checks were made which included a planned TRV event. The stainless steel drum-sampling event also occurred Wednesday.

There was an unplanned TRV event Friday morning, April 16, 1999 due to a brief power outage.

Received bottom ash analytical reports for the thermally treated soils processed from Thursday, April 8, 1999 to Saturday, April 17, 1999. The analytical lab's test results indicated that the thermally treated material met the project's organic criteria for backfill conditions.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood management, feed preparation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities, the weekly OSHA method 93 test and the weekly NIOSH 5518 tests for BNA.

On Thursday, April 22, 1999, OHM completed incineration phase of the on site soils. Much of the materials excavated consisted of fill soils and the naturally occurring soils at the bottom of the excavation. Minor amounts of debris and wood are left outside the feed preparation building to be sent for offsite disposal. Removal of HDPE liner and protective filter fabric will commence once the sprung structure fabric has been removed.

During the time frame of this report, 4 roll-off containers with various amounts of PPE, lime bags, wood and miscellaneous site metal were shipped off-site as non-hazardous waste to a hazardous waste landfill.

Related site activities included a weekly progress meeting held on Tuesday, April 20, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance. A Senior Partnering Meeting was held on Wednesday, April 21, 1999. Those present included Mr. Tom Voltaggio, Mr. Gregg Crystall of EPA; Mr. Bill Parsons, Mr. Mike Welsh, Mr. Larry Newcomer, Ms. Sandy Downs of PADEP; Mr. Lloyd Caldwell, Mr. Denis duBreuil, Mr.Jared Olsen of USACE; Mr. Matt Radek of OHM.

Major activities for the next week includes start of site demolition, removal of steam scrubber system, acid gas scrubber purging, ice removal from the O2 pad, removal of oxygen tanks and system, WWTP operations, ash handling, final ash backfilling, debris handling and the implementation of the project close out plan.

The incineration phase of the project is tentatively scheduled to be complete April 24, 1999. When the analytical test results from the final feed soils are indicated to meet the site backfill criteria. The incinerator will then be turned off. Demobilization of the unit is scheduled to begin once the unit sufficiently cools.

PREPARED BY: John Ricard for Michael Ogden (USACE)


Wednesday, April 14, 1999

FOR DATES: From 7:00 a.m. Wednesday, Apr. 7, 1999 to 7:00 a.m. Wednesday, Apr. 14, 1999.

The Thermal Destruction Facility (TDF) was on-line for 154.8 hours and processed 4927.2 tons of contaminated soil during the time frame of this report. The TDF was on-line 92.1 % of the time with an average hourly feed rate of 31.8 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 284,714.7 tons of contaminated soil, which amounts to approximately 98.2% of the overall contaminated soil on-site.

Hazardous waste feed was manually stopped on Thursday, April 8, 1999 in order to investigate problems with the quench tower spray nozzles. The nozzles were cleaned, reassembled and reinstalled. Feed was off for approximately 10 hours. Hazardous waste feed was manually stopped on Friday, April 9, 1999 during a lightning storm as part of the standard operating procedure to prevent power loss to the TDF. Feed resumed once the storm had passed.

Received bottom ash analytical reports for the thermally treated soils processed from Friday, April 3, 1999 to Saturday, April 4, 1999. The analytical lab's test results indicated that the thermally treated soil stored in bin 6 did not meet the project's organic criteria due to the concentration of the semi-volatile organic compound Beta-Naphthylamine being greater than the organic criteria for backfilling the thermally treated soil on-site. The contents of bin 6 will be returned to the exclusion zone for further thermal treatment. Fly ash bin E was taken down so the contractor can move the failed bottom ash to the exclusion zone for further thermal treatment. The contents of bin 4, failing bottom ash that had been produced on April 2nd and 3rd, was returned to the exclusion zone with the contents of bin 6.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood management, feed preparation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities, the weekly OSHA method 93 test and the weekly NIOSH 5518 tests for BNA.

On Monday, April 12, 1999, OHM completed the excavation of the unlined portion of the exclusion zone. Much of the materials excavated consisted of fill soils and the naturally occurring soils at the bottom of the excavation. Minor amounts of demolition debris and wood were encountered. Once the materials were excavated and the floor of the excavation was confirmed to be at elevation 545 feet msl, shot rock and screened site rock was backfilled against the sheet piling. The remaining soils to be thermally treated are wholly contained on the HDPE liner adjacent to and inside of the feed preparation building. These soils will be removed with the mini-excavator and hand held equipment in order to maintain the integrity of the liner.

During the time frame of this report, 5 roll-offs containing various amounts of PPE, lime bags, wood and miscellaneous site metal were shipped off-site as non-hazardous waste to a hazardous waste landfill.

Related site activities included a weekly progress meeting held on Tuesday, April 13, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, rock backfilling, and the implementation of the project close out plan.

The incineration phase of the project is tentatively scheduled to be complete during the week of April 18, 1999. When the analytical test results from the final feed soils are indicated to meet the site backfill criteria, the incinerator will be turned off. Demobilization of the unit is scheduled to begin once the unit sufficiently cools.

PREPARED BY: Steve Gibson for Frank Cleaver (USACE)


Wednesday, April 7, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 31, 1999 to 7:00 a.m. Wednesday, April 7, 1999.

The Thermal Destruction Facility (TDF) was on-line for 164.2 hours and processed 5858.2 tons of contaminated soil during the time frame of this report. The TDF was on-line 97.7 % of the time with an average hourly feed rate of 35.7 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 281,588 tons of contaminated soil, which amounts to approximately 97.7% of the overall contaminated soil on-site.

During the time frame of this report, hazardous waste feed was manually stopped for a brief period of time on Saturday, April 3, 1999 when one of the power grids in the feed preparation building broke down while the back-up power grid was being repaired. Feed resumed 2 hours later, once one of the power grids was operational.

Received bottom ash analytical reports for the thermally treated soils processed from Friday, April 2, 1999 to Saturday, April 3, 1999. The analytical lab's test results indicated that the thermally treated soil stored in bin 4 did not meet the project's organic criteria due to the concentration of the semi-volatile organic compound Beta-Naphthylamine being greater than the organic criteria for backfilling the thermally treated soil on-site. The contents of bin 4 will be returned to the exclusion zone for further thermal treatment. Fly ash bin E is currently being taken down so the contractor can move the ash from the ash pad to the exclusion zone, which has been significantly reduced as the thermal phase of the project comes to a close.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood management, feed preparation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. ConCover was placed on the sides of the ash stockpile to help prevent erosion and fugitive emissions. Shot rock was imported from off-site to backfill the excavation floor for the 2-foot rock bottom. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation activities were continued south of the feed preparation building, adjacent to the CZ113 sheet piling. Much of the materials excavated consisted of fill soils and the naturally occurring soils at the bottom of the excavation. Some minor undercutting was performed when fill material was observed at elevation 545. Naturally occurring soil was encountered approximately 6-inches below the normal excavation floor elevation. No drums or demolition debris were encountered. Once the materials were excavated, shot rock and screened site rock was backfilled against the sheet piling. The contractor continued to remove the soil on top of the HDPE liner outside of the feed preparation building on the northern side of the CZ113 sheet piling. The soils were removed with a mini-excavator and hand held equipment in order to maintain the integrity of the liner.

During the time frame of this report, 4 roll-offs containing various amounts of PPE, lime bags, wood and miscellaneous site metal were shipped off-site as non-hazardous waste to a hazardous waste landfill. The roll-off containing the materials bulked from the drums excavated on-site was shipped off-site on Tuesday, April 6, 1999. The analytical test results for the drum bulk composite indicated that the contents of the roll-off were non-hazardous. The roll-off was shipped off-site as non-RCRA, non-DOT regulated waste to a hazardous waste landfill.

Related site activities included:

* A weekly progress meeting was held on Tuesday, April 6, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

* The fourth and final stack testing event for metals and particulate was completed on Wednesday, March 31, 1999.

* Kurt Davis, Technical Assistance Grant Representative, was on-site to observe stack testing.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, rock backfilling, and the implementation of the project close out plan.

PREPARED BY: Steve Gibson for William Werntges (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.